### v1

# Chiral Domain Walls Induced by Radially Magnetized Nanotube Geometry

Abstract
We theoretically investigate chiral domain walls (DWs) formed in radially magnetized nanotubes with perpendicular magnetic anisotropy (PMA). Unlike tubes with an easy axis of magnetization oriented in other directions, radially magnetized tubes exhibit geometrical effects arising not only from exchange interactions but also from magnetostatic interactions, which lead to Dzyaloshinskii-Moriya interaction (DMI)-like effects. We derive expressions for the effective magnetic fields acting on DWs within PMA nanotubes and quantify SOT-driven DW motion using an analytical one-dimensional model, which is validated by micromagnetic simulations. Our results show that the DMI-like field due to magnetostatic interactions can be as significantly as the contribution of the material-induced DMI in nanotubes with diameters below 100 nm. This implies that the direction and speed of DW motion in PMA nanotubes could differ from those observed in nanoribbons composed of the same material. Furthermore, we demonstrate that DW velocity can be effectively controlled by adjusting the tube diameter and exchange stiffness constant of the magnetic layer, rather than relying solely on the material-induced DMI. These insights are expected to greatly expand the potential applications of PMA nanotube-based DW devices.
I Introduction
Recent studies on the behavior of magnetic textures in curved magnetic layers have garnered significant attention [1, 2, 3]. It has been theoretically predicted that the curvature and torsion of magnetic films affect the dynamics of magnetic domain walls (DWs) [4, 5, 6, 7], with some predictions corroborated by experimental evidence [8, 9, 10, 11, 12]. Understanding these phenomena could enable the storage and manipulation of DWs within three-dimensional (3D) structures [3, 13], paving the way for high-density data storage applications [14, 15].
In this study, we focus on nanotubes as a promising form for future 3D magnetic devices, alongside nanoribbons and nanocylinders. While fabricating 3D nanoribbon structures requires advanced techniques to deposit magnetic layers onto sidewalls [16], nanotubes and nanocylinders could be more readily fabricated using existing methods such as electroplating [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30] and atomic layer deposition [31, 32]. Moreover, nanotube structures can achieve a smaller cross-sectional area compared to non-hollow nanocylinders, thereby reducing the shift current for DWs.
Previous theoretical research on the properties of DWs in nanotubes has primarily focused on systems with in-plane magnetized anisotropy (IMA) [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44]. These studies have predominantly considered DW motion driven by conventional spin-transfer torque (STT) and Ørsted field [37, 44]. In contrast, this study aims to establish a theoretical framework for DW dynamics in radially magnetized nanotubes induced by perpendicular magnetized anisotropy (PMA) materials [45]. Such systems are expected to utilize spin-orbit torque (SOT) [46, 47, 48], potentially enabling more efficient DW motion compared to traditional driving forces. Therefore, PMA nanotube-based DW shift registers are considered highly promising candidates for low power consumption data storage applications.
Generally, geometrical effects are known to arise from exchange interactions [49, 50]; however, it has been predicted that geometrical effects due to magnetostatic interactions can also occur, depending on the curvature characteristics and the easy axis of magnetization [51]. Previous research indicates that such magnetostatic interactions induce geometrical effects only in systems with a radial easy axis of magnetization in tube structures, leading to energy contributions similar to Dzyaloshinskii-Moriya interaction (DMI) [52, 53]. This suggests the potential induction of chirality in DWs within tube structures. The magnitude of DMI is crucial for the efficiency of SOT-driven DW motion [46], making it essential to theoretically predict the impact of these geometrical effects, to evaluate the potential applications of PMA nanotubes.
The purpose of this study is to elucidate the characteristics of chiral DWs in PMA nanotubes and to quantitatively predict the influence of geometrical effects on SOT-driven DW motion. Specifically, we derive expressions for the effective magnetic fields acting on DWs due to magnetostatic interactions and formulate an analytical model for SOT-driven DW motion. The validity of this model is confirmed through micromagnetic (µM) simulations.
The remainder of this paper is organized as follows. In Section II, we describe the model of PMA nanotubes. Section III presents the expressions for effective magnetic fields due to geometrical effects in PMA nanotubes and discusses the characteristics of these fields and the stability of DWs as a function of nanotube diameter. In Section IV, we present the results of SOT-driven DW motion calculations, comparing the outcomes from the analytical one-dimensional model (1DM) [54, 55, 46, 56, 57] and µM simulations. We also discuss the comparative shift characteristics of PMA nanotubes and nanoribbons, as well as the impact of variations in tube diameter and magnetic properties on these characteristics. Finally, Section V summarizes the key findings and implications of this study.
II Model of radially magnetized nanotube
We consider a magnetic nanotube that extends in the $z$ direction, with inner and outer radii denoted as $R_{0}$ and $R_{1}$, respectively (Fig. 1). The thickness of the nanotube is defined as $t=R_{1}-R_{0}$, with the assumption that $t\ll R_{0}$. We assume that the nanotube is of infinite length, allowing us to neglect any end effects. Our study focuses on DWs in nanotubes composed of heavy metal (HM)/ferromagnet (FM) bilayers, where the interfacial DMI induces SOT-driven DW motion [58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69]. The normalized magnetization vector is expressed using the basis vectors of the cylindrical coordinate system: $\bm{m}=m_{r}\bm{e}_{r}+m_{\varphi}\bm{e}_{\varphi}+m_{z}\bm{e}_{z}$, where $\bm{e}_{r}$, $\bm{e}_{\varphi}$, and $\bm{e}_{z}$ are the radial, azimuthal, and axial directions, respectively.
Figure 1: Schematic illustration of a radially magnetized nanotube containing a DW. The magnetic structure for a topological charge of $Q=+1$ is presented. The enlarged diagram shows the magnetization configuration of the DW as seen from the inside of the tube.
[…]
For PMA layers of nanotubes, the effective magnetic anisotropy constant is approximately given by
$K_{{\rm eff}}=K_{{\rm u}}-2\pi M_{{\rm s}}^{2}-A/[R_{1}(R_{1}-t/2)],$ (1)
where the second term represents the demagnetization energy density, and the third term accounts for the exchange energy density associated with the radial domain state [70], which is specific to tube structures. Figure 2 shows the phase diagram of magnetic domain structures in the PMA nanotube, validated through µM simulations (see Appendix A). When $K_{{\rm eff}}>0$, the easy axes of magnetization are oriented radially; otherwise, the easy axis aligns along the axial direction.
Figure 2: Phase diagram of magnetic domain structures in the PMA nanotube. The dashed line represents the boundary of stable states as determined by Eq. (1). The markers indicate conditions validated through µM simulations. The following parameters are used; $A=1\,{\rm uerg/cm}$ and $t=1\,{\rm nm}$.
In PMA nanotubes, the energy densities associated with the PMA, DMI and exchange interactions can be expressed respectively as $\varepsilon_{{\rm PMA}}=-K_{{\rm eff}}m_{r}^{2}$, $\varepsilon_{{\rm DMI}}=D\left(m_{r}\frac{\partial}{\partial z}m_{z}-m_{z}\frac{\partial}{\partial z}m_{r}\right)$, and $\varepsilon_{{\rm exc}}=A\sum_{i=x,y,z}\left[\left(\frac{\partial m_{i}}{\partial r}\right)^{2}+\frac{1}{r^{2}}\left(\frac{\partial m_{i}}{\partial\varphi}\right)^{2}+\left(\frac{\partial m_{i}}{\partial z}\right)^{2}\right]$, where $m_{x}=m_{r}\cos\varphi-m_{\varphi}\sin\varphi$ and $m_{y}=m_{r}\sin\varphi+m_{\varphi}\cos\varphi$. The energy of magnetistatic interaction is expressed as:
$E_{{\rm dem}}$ $=E_{{\rm dem,bb}}+E_{{\rm dem,ss}}+E_{{\rm dem,bs}},$ (2)
$E_{{\rm dem,bb}}$ $=\frac{1}{2}\int{\rm d}^{3}\bm{x}\int{\rm d}^{3}\bm{x}^{\prime}\frac{\rho_{{\rm m}}\left(\bm{x}\right)\rho_{{\rm m}}\left(\bm{x}^{\prime}\right)}{\left|\bm{x}-\bm{x}^{\prime}\right|},$
$E_{{\rm dem,ss}}$ $=\frac{1}{2}\int{\rm d}^{2}\bm{x}\int{\rm d}^{2}\bm{x}^{\prime}\frac{\sigma_{{\rm m}}\left(\bm{x}\right)\sigma_{{\rm m}}\left(\bm{x}^{\prime}\right)}{\left|\bm{x}-\bm{x}^{\prime}\right|},$
$E_{{\rm dem,bs}}$ $=\int{\rm d}^{3}\bm{x}\int{\rm d}^{2}\bm{x}^{\prime}\frac{\rho_{{\rm m}}\left(\bm{x}\right)\sigma_{{\rm m}}\left(\bm{x}^{\prime}\right)}{\left|\bm{x}-\bm{x}^{\prime}\right|},$
where $\sigma_{{\rm m}}\left(\bm{x}\right)=M_{\rm s}\bm{m}\left(\bm{x}\right)\cdot\bm{n}_{\perp}\left(\bm{x}\right)$ is the surface magnetic charge density, with $\bm{n}_{\perp}\left(\bm{x}\right)$ being the normal vector of the tube surface, and $\rho_{{\rm m}}\left(\bm{x}\right)=-M_{\rm s}\nabla\cdot\bm{m}\left(\bm{x}\right)$ is the volume magnetic charge density [70].
III Characteristics of domain walls
In this section, we derive the energy densities per nanotube cross section and present expressions for the effective fields acting on the DWs in PMA nanotubes.
[…]
The PMA, DMI, and exchange energy densities per nanotube cross section are expressed as $w_{{\rm PMA,DMI,exc}}=\int{\rm d}z\,\varepsilon_{{\rm PMA,DMI,exc}}$. The integral result of the PMA term is $w_{{\rm PMA}}=2\sqrt{AK_{{\rm eff}}}$, which does not depend on $\phi$ and therefore does not affect the DW dynamics (see Eq. (9)).
The integral result of the DMI term is given by
$w_{{\rm DMI}}\left(\phi\right)=-\pi\lambda M_{{\rm s}}H_{{\rm DMI}}Q\cos\phi,\$ (3)
where $H_{{\rm DMI}}=\frac{D}{M_{{\rm s}}\lambda}$ is the material-induced DMI field [71, 56, 72].
The integral result of the $\phi$-dependent exchange interaction term is given by
$w_{{\rm exc}}\left(\phi\right)=\lambda M_{{\rm s}}H_{{\rm a,exc}}^{{\rm tube}}\cos^{2}\phi,\$ (4)
where $H_{{\rm a,exc}}^{{\rm tube}}=-\frac{4A}{M_{{\rm s}}}\frac{1}{R_{1}^{2}-R_{0}^{2}}\ln\frac{R_{1}}{R_{0}}$ is the curvilinear-geometry-induced anisotropy-like field [49, 50]. This result is obtained by integrating $\varepsilon_{{\rm exc}}=\frac{A}{r^{2}}\left(1+\frac{r^{2}}{\lambda^{2}}{\rm sech}^{2}\frac{z-X}{\lambda}-{\rm sech}^{2}\frac{z-X}{\lambda}\cos^{2}\phi\right)$ with respect to $X$ and is consistent with the result for IMA tubes with an azimuthal easy axis of magnetization [44]. In the case where $H_{{\rm a,exc}}^{{\rm tube}}$ is positive (negative), the Bloch (Néel)-type DW is favored.
The energy density of the $\phi$-dependent magnetostatic interaction term is given by the following equation, as derived in Appendix B:
$w_{{\rm dem}}\left(\phi\right)=\lambda M_{{\rm s}}\left(H_{{\rm a,dem}}^{{\rm{\rm tube}}}\cos^{2}\phi-\pi H_{{\rm D,dem}}^{{\rm tube}}Q\cos\phi\right),$ (5)
where $H_{{\rm a,dem}}^{{\rm{\rm tube}}}$ and $H_{{\rm D,dem}}^{{\rm tube}}$ represent the anisotropy-like demagnetizing field within the DW [73] and DMI-like field which is specific to radially magnetized tube [51], respectively. These fields are expressed as
$H_{{\rm a,dem}}^{{\rm{\rm tube}}}=\frac{4\pi M_{{\rm s}}}{\rho_{1}^{2}-\rho_{0}^{2}}\int_{-\infty}^{+\infty}{\rm d}x\int_{0}^{+\infty}{\rm d}y\,\frac{\tanh x\left(\rho_{1}^{2}I_{11}^{-1}\left(\rho_{1},\rho_{1},y\right)-2\rho_{1}\rho_{0}I_{11}^{-1}\left(\rho_{1},\rho_{0},y\right)+\rho_{0}^{2}I_{11}^{-1}\left(\rho_{0},\rho_{0},y\right)\right)}{\cosh x\cosh\left(y-x\right)}\,,$ (6)
$H_{{\rm D,dem}}^{{\rm tube}}=\frac{8M_{{\rm s}}}{\rho_{1}^{2}-\rho_{0}^{2}}\int_{-\infty}^{+\infty}{\rm d}x\int_{0}^{+\infty}{\rm d}y\int_{\rho_{0}}^{\rho_{1}}{\rm d}\rho\,\frac{\tanh x\left(\rho_{1}I_{11}^{1}\left(\rho_{1},\rho,y\right)-\rho_{0}I_{11}^{1}\left(\rho_{0},\rho,y\right)\right)\rho}{\cosh\left(y-x\right)},$ (7)
where $\rho_{0,1}=R_{0,1}/\lambda$, $I_{\alpha\beta}^{\mu}\left(a,b,s\right)=\int_{0}^{\infty}{\rm d}x\,x^{\mu}J_{\alpha}\left(ax\right)J_{\beta}\left(bx\right){\rm e}^{-sx}$ and $J_{n}\left(x\right)$ is the $n$-th order Bessel function of the first kind. The integrals are provided more specifically as follows [74]: $I_{11}^{1}\left(a,b,s\right)=\frac{s\kappa}{2\pi ab\sqrt{ab}}\left(\frac{1-\frac{\kappa^{2}}{2}}{1-\kappa^{2}}E\left(\kappa\right)-K\left(\kappa\right)\right)$ and $I_{11}^{-1}\left(a,b,s\right)=\frac{s}{\pi\sqrt{ab}}\left[\frac{1}{\kappa}E\left(\kappa\right)-\frac{\kappa\left(a^{2}+b^{2}+\frac{s^{2}}{2}\right)}{2ab}K\left(\kappa\right)\right]+\frac{1}{4ab}\left[a^{2}+b^{2}+\left(a^{2}-b^{2}\right){\rm sgn}\left(a-b\right)\left(\Lambda\left(\nu,\kappa\right)-1\right)\right]$, where $E\left(k\right)=\,\int_{0}^{\pi/2}{\rm d}\theta\sqrt{1-k^{2}\sin^{2}\theta}$ is the first elliptic integral function, $K\left(k\right)=\,\int_{0}^{\pi/2}{\rm d}\theta\left(\sqrt{1-k^{2}\sin^{2}\theta}\right)^{-1}$ is the second elliptic integral function, $\Pi\left(n,k\right)=\,\int_{0}^{\pi/2}{\rm d}\theta\left[\left(1-n\sin^{2}\theta\right)\sqrt{1-k^{2}\sin^{2}\theta}\right]^{-1}$ is the third elliptic integral function, $\kappa\left(a,b,s\right)=2\sqrt{\frac{ab}{\left(a+b\right)^{2}+s^{2}}}$, $\nu\left(a,b\right)=\frac{4ab}{\left(a^{2}+b^{2}\right)}$, and $\Lambda\left(n,k\right)=\frac{2}{\pi}\sqrt{\left(1-n\right)\left(n-k^{2}\right)/n}\,\Pi\left(n,k\right)$. As shown in Fig. 3, the results of energy densities obtained from our µM simulations closely align with those derived from the equations presented in Eqs. (4), (5), (6), and (7).
Figure 3: Energy densities per cross section in PMA nanotubes as a function of magnetization angle. The following parameters are used; $A=1\,{\rm uerg/cm}$ and $K_{\rm u}=8.3\,{\rm Merg/cm}^{3}$.
The underlying physical mechanisms of the two terms in Eq. (5) are shown in Fig. 4. The first term of Eq. (5) arises from the magnetic charges in the DW (Fig. 4(a)). From the perspective of the unified theoretical framework [51], this phenomenon can be classified as the interactions between geometrical magnetostatic charges. Since the DW in the radially magnetized tube does not generate a magnetic charge in the perfect Bloch-type DW ($\phi=\pm\pi/2$), these states are minimal for the term.
The second term of Eq. (5) arises from the stray field from the magnetic domains in the radially magnetized tube (Fig. 4(b)). When the DW magnetization is aligned with the stray field, corresponding to the right-handed Néel DW, the energy of this configuration is at a minimum. This term consists of not only the interaction between geometrical magnetostatic charges but also the interaction between geometrical and surface magnetostatic charges [51]. The fundamental basis of this field lies in the asymmetry in the distribution of magnetic charges on the front and back surfaces of the magnetic layer. It should be noted that this field is specific to radially magnetized tubes and is not present in ribbons and other tubes with a different easy axis of magnetization.
Figure 4: Schematic illustration of geometrical effects arising from magnetostatic interactions in PMA tubes containing a DW.
Figure 5 shows the radius dependence of the each effective field acting on the DW in the PMA nanotube, as evaluated using the derived equations. The absolute values of the effective fields $H_{{\rm a,exc}}^{{\rm tube}}$ and $H_{{\rm D,dem}}^{{\rm tube}}$ increase as the tube radius decreases, while $H_{{\rm a,dem}}^{{\rm tube}}$ remains approximately constant regardless of the radius.
The exchange interaction increases with greater curvature, leading to a corresponding increase in $\left|H_{{\rm a,exc}}^{{\rm tube}}\right|$ as the radius decreases, shown in Fig. 5(a). Since this field is negative, Néel-type DWs are favored in systems with smaller radii. This characteristics is similar to the behavior observed in nanoribbons, where narrower wire width favor Néel-type DWs [73]. However, unlike in nanoribbons where this characteristics is attributed to exchange interactions, in nanotube, it is due to magnetostatic interactions, indicating that the origins of the two characteristics are different.
Figure 5: Radius dependence of the effective field acting on the DW in PMA nanotubes with $t=1\,{\rm nm}$.
The anisotropy-like demagnetization field from the magnetic charges in the DW, $H_{{\rm a,dem}}^{{\rm tube}}$, is primarily determined by the DW width and is thus independent of the tube radius, as shown in Fig. 5(b). This characteristics differs from the behavior observed in nanoribbons, where the anisotropy-like demagnetization field depends on the wire width [73]. In nanoribbons, as the wire width decreases, this effective field decreases and eventually reverses its sign at a certain width, favoring Néel-type DWs over Bloch-type DWs. This discrepancy arises from the presence of edges in nanowires. The absence of edges in tubes prevents the emergence of magnetic charges within Bloch-type DWs, making them preferred ($H_{{\rm a,dem}}^{{\rm tube}}>0$).
The stray field generated by the magnetic domains, $H_{{\rm D,dem}}^{{\rm tube}}$, decreases as the radius increases, as shown in Fig. 5(c). This is because the distance between opposing magnetic domains located on either side of the tube’s central axis increases. Furthermore, a reduction in the DW width corresponds to the expansion of the magnetic domain regions, thereby enhancing the stray field.
In the radially magnetized tube, the total anisotropy and DMI fields acting on a DW are written by $H_{{\rm a}}=H_{{\rm a,exc}}^{{\rm tube}}+H_{{\rm a,dem}}^{{\rm tube}}$ and $H_{{\rm D}}=H_{{\rm D,dem}}^{{\rm tube}}+H_{{\rm DMI}}$, respectively. The magnetization angle of the DW in the equilibrium states can be expressed by the following equation [75]:
$\phi_{{\rm eq}}=\begin{cases}0&\frac{2H_{{\rm a}}}{\left|\pi H_{{\rm D}}\right|}<1\cap H_{{\rm D}}>0\\ \pi&\frac{2H_{{\rm a}}}{\left|\pi H_{{\rm D}}\right|}<1\cap H_{{\rm D}}<0\\ {\rm arccos}\left(\frac{\pi H_{{\rm D}}}{2H_{{\rm a}}}\right)&\frac{2H_{{\rm a}}}{\left|\pi H_{{\rm D}}\right|}>1\end{cases},$ (8)
[…]
In nanoribbons, $H_{{\rm a}}$ arises exclusively from the anisotropy-like demagnetization field arising from magnetic charges within the DW, while $H_{{\rm D}}$ does not include any geometry-dependent terms. Consequently, in nanoribbons, the total anisotropy field is influenced by the nanoribbon width [73], whereas the total DMI field remains invariant to changes in dimensions.
Figure 6: Magnetization angle of DW in equilibrium states of PMA nanotubes. The following parameters are used; $A=1\,{\rm uerg/cm}$, $K_{\rm u}=8.3\,{\rm Merg/cm}^{3}$, and $t=1\,{\rm nm}$.
IV Spin-orbit torque-driven domain wall motion
In this section, we present the numerical results of current-induced DW motion in PMA nanotubes under the influence of STT and SOT.
The dynamics of the DW are described by the following differential equations within the framework of an analytical 1DM [54, 55, 46, 56, 57]:
$\frac{{\rm d}\phi}{{\rm d}t}$ $=\frac{\gamma}{1+\alpha^{2}}\left(-Q\Gamma_{1}-\alpha\Gamma_{2}\right),$ (9)
$\frac{{\rm d}\left(X/\lambda\right)}{{\rm d}t}$ $=\frac{\gamma}{1+\alpha^{2}}\left(-\alpha\Gamma_{1}+Q\Gamma_{2}\right),$
$\Gamma_{1}$ $=Q\frac{\pi}{2}\xi_{{\rm DLT}}j\cos\phi,$
$\Gamma_{2}$ $=Q\xi_{{\rm STT}}j+\frac{1}{2M_{{\rm s}}\lambda}\frac{\partial w}{\partial\phi},$
where $\gamma$ is the gyromagnetic ratio, $\alpha$ is the Gilbert damping, $j$ is the current density, and $w=w_{{\rm DMI}}\left(\phi\right)+w_{{\rm exc}}\left(\phi\right)+w_{{\rm dem}}\left(\phi\right)$ is the energy density per cross-sectional area of the nanotube. The damping-like torque (DLT) efficiency induced by spin Hall effects is given by $\xi_{{\rm DLT}}=\frac{\hbar\theta_{{\rm SH}}}{2M_{{\rm s}}\left|e\right|t}$, and the STT efficiency is defined as $\xi_{{\rm STT}}=-\frac{\hbar P}{2M_{{\rm s}}\left|e\right|\lambda}$. Here, $e$ is the elementary charge, $\hbar$ is the Dirac constant, $\theta_{{\rm SH}}$ is the spin Hall angle (SHA) , and $P$ is the spin polarization. We assume that the same current density is present in both the HM and FM layers [69]. Based on the findings from the previous section, $\Gamma_{2}$ in Eq. (9) can be expressed as
$\Gamma_{2}=$ $\,\xi_{{\rm STT}}j-\frac{1}{2}\left(H_{{\rm a,exc}}^{{\rm tube}}+H_{{\rm a,dem}}^{{\rm tube}}\right)\sin 2\phi$ (10)
$+Q\frac{\pi}{2}\left(H_{{\rm D,dem}}^{{\rm tube}}+H_{{\rm DMI}}\right)\cos\phi.$
We assume that the HM layer consists of materials with a negative SHA, such as W, Ta, or Hf [62, 64, 76, 69]. In the case where the FM layer is composed of CoFeB, it is known that the sign of DMI is positive for a W layer [77, 69], negative for a Hf layer [64], and can take both signs for a Ta layer [62, 64, 78]. Consequently, we carried out simulations for different values of DMI: positive, negative, and zero.
We disregard the field-like torque (FLT) as it does not serve as a driving force for DW motion in PMA layers. The FLT acts on the DW as an effective field directed along the azimuthal axis in PMA nanotubes. For example, it has been reported that in HM/CoFeB/MgO structures with HM layers composed W or Ta, the FLT is approximately one order of magnitude smaller than the DLT [79]. Although the contribution of the Ørsted field is included in the FLT, for the parameters of interest (specifically, $R_{1}=O({10\,\rm nm})$ and $t=O({1\,\rm nm})$), this contribution is more than an order of magnitude smaller than the DLT [80].
Figure 7 shows the transient position of the DW in the PMA nanotube under applied currents. The results from 1DM and µM simulations are consistent, demonstrating the validity of the approximation that the DW width is constant in the 1DM. The magnetization angle of the DW remains constant during current flow, and the direction of DW motion depends on the sign of $D$. The DWs driven by SOT are slightly displaced from the perfect Bloch-type DW ($\phi=\pi/2$). In systems with a negative SHA, it has been theoretically shown that the DW moves in the direction of current (electron) flow for the right (left)-handed Néel-type DW [81, 69]. Our computational results are consistent with this prediction. These results confirm that the SOT-driven motion in PMA nanotubes is qualitatively similar to that observed in PMA nanoribbons.
Figure 7: Transient position of DW in the PMA nanotube under applied currents. Images are snapshots of the magnetization structure by µM simulations. The enlarged views show the magnetization structure in the vicinity of the DW, with the green arrows indicating the direction of magnetization of the DW. The following parameters are used; $A=1\,{\rm uerg/cm}$, $K_{\rm u}=8.3\,{\rm Merg/cm}^{3}$, $\alpha=0.2$, $P=0.5$, $\theta_{{\rm SH}}=-0.2$, $R_{1}=25\,{\rm nm}$, and $t=1\,{\rm nm}$.
Figure 8 shows the DW velocity as a function of the current density ($v$-$j$ curve) for the PMA nanotube. To clarify the geometrical effects, we also present the results of PMA nanoribbon. In systems where only SOT is present without STT, a comparison of the $v$-$j$ curves for the nanotube and nanoribbon ( Fig. 8(a) and (b)) reveals a more pronounced influence of the unique DMI-like field in the PMA nanotube. In the nanoribbon, the direction of DW motion reverses with a change in the sign of $D$, but the speeds of the DW motion are the same. In contrast, in the nanotube, the direction of DW motion does not necessarily reverse with a change in the sign of $D$, and the speed of DW motion varies accordingly. It should be noted that geometry-induced positive DMI contribution is always present in our PMA nanotube models.
Figure 8: DW velocity as a function of the current density ($v$-$j$ curve) in the PMA nanotube with $R_{1}=25\,{\rm nm}$ and the PMA nanoribbon with a width of $50\,{\rm nm}$. The anisotropy-like field ($H_{\rm a}$) used for the nanoribbons in the 1D simulation was derived from energy calculations obtained through µM simulations. The pulse widths for the calculations are 2 ns. The following parameters are used; $A=1\,{\rm uerg/cm}$, $K_{\rm u}=8.3\,{\rm Merg/cm}^{3}$, $\alpha=0.2$, $\theta_{{\rm SH}}=-0.2$, and $t=1\,{\rm nm}$.
In systems where both STT and SOT are present, SOT serves as the driving force for DW motion, while STT indirectly affects DW motion by modulating the DW magnetization angle [69]. In the case of $D=+0.1\,{\rm erg/cm^{2}}$, the higher DW speed observed in the nanotube (Fig. 8(c)) compared to the nanoribbon (Fig. 8(d)) in the low current density region. This is attributed to the presence of a specific positive-sign DMI-like field within the PMA nanotube. The decrease in DW velocity with increasing current density in the high current region is the characteristic of systems with a negative SHA and a positive DMI [69]. This behavior reflects the transition of the DW from a right-handed Néel-type DW to a left-handed Néel-type DW, driven by increasing STT with higher currents.
In PMA nanotubes, a geometry-induced positive DMI-like field enables DW motion in the direction of the current flow even when $D\leq 0$ and $\theta_{\rm SH}<0$, which differs from PMA nanoribbons [81]. As shown in Fig. 8(d), the DW in the PMA nanoribbon only moves in the direction of electron flow when $D\leq 0$, whereas in the PMA nanotube, the DW moves in the direction of the current flow in the low current density region even when $D\leq 0$ shown in Fig. 8(c). The behavior observed in the PMA nanotube is consistent with the total DMI field, where $H_{{\rm D}}=H_{{\rm D,dem}}^{{\rm tube}}+H_{{\rm DMI}}>0$.
In PMA nanotubes, even when the material-induced DMI is weak, geometrical effects can stabilize a Néel-type DW structure during current flow, enabling higher speed motion. Figure 9 shows the $v$-$j$ curve for PMA nanotube with $D=0$ for different combinations of nanotube radius and exchange stiffness of the magnetic layer. As the radius decreases, both $H_{{\rm D,dem}}^{{\rm tube}}$ and $\left|H_{{\rm a,exc}}^{{\rm tube}}\right|$ increase (Fig. 5). This leads to the stabilization of a right-handed Néel-type DW (Fig. 6), which in turn increases the speed of DW motion. Additionally, an increase in exchange stiffness results in a larger $\left|H_{{\rm a,exc}}^{{\rm tube}}\right|$, further enhancing the stabilization of the Néel-type DW.
Figure 9: Comparison of $v$-$j$ curves for PMA nanotubes with different radius $(R_{1})$ and exchange stiffness constant $(A)$. The pulse widths for the calculations are 5 ns. The following parameters are used; $K_{\rm u}=8.3\,{\rm Merg/cm}^{3}$, $D=0$, $\alpha=0.2$, $P=0.5$, $\theta_{{\rm SH}}=-0.2$, and $t=1\,{\rm nm}$.
The velocity of DW motion induced by the SOT for the Néel-type DW in the low current regime is described by the following equation [57]:
$v=-\frac{\gamma\lambda}{\alpha}\frac{\pi}{2}\xi_{\rm DLT}j\left[1-\frac{\left(\alpha^{-1}\frac{\pi}{2}\xi_{\rm DLT}+\xi_{\rm STT}\right)^{2}}{2\left(\frac{\pi}{2}H_{\rm D}-H_{\rm a}\right)^{2}}j^{2}\right].$ (11)
The second term reflects the deceleration that occurs when the DW deviates from the perfect Néel-type DW ($\phi=0,\pi$) during current flow. As the radius of the nanotube increases or the exchange stiffness increases, the denominator of this second term becomes larger (note that $H_{\rm D}>0$ and $H_{\rm a}<0$), resulting in a reduction of the deceleration effect. Consequently, the speed of the DW increases.
V Conclusions
In this study, we analytically investigated the characteristics of chiral DWs induced by geometrical effects in radially magnetized nanotubes with PMA. We derived expressions for the effective magnetic fields acting on DWs and formulated the SOT-driven DW motion in PMA nanotubes using a one-dimensional theory. This theory was validated through micromagnetic simulations.
Our findings demonstrate that the stray field originating from the magnetostatic interactions specific to PMA nanotubes can have a contribution comparable to the material-induced DMI for diameters in the sub-$100\,{\rm nm}$ range. Consequently, the direction and speed of DW motion in PMA nanotubes may differ from those observed in nanoribbons composed of the same material.
The DMI-like field resulting from magnetostatic interactions depends on the tube diameter, enabling the control of SOT-driven DW motion speeds through geometric adjustments rather than the material-induced DMI, which is sensitive to deposition conditions [82, 77, 69]. Additionally, the anisotropy-like field induced by curvilinear geometry, arising from exchange interactions, helps to stabilize Néel-type DWs as tube diameters decrease or exchange stiffness constants of the magnetic layer increase. Utilizing these effects is expected to result in more efficient SOT-driven DW motion. These insights are anticipated to significantly expand the potential applications of PMA nanotubes.
Appendix A Micromagnetic simulations
We conducted micromagnetic (µM) simulations of a DW inside a PMA nanotube to validate our analytical formulation. Our µM program utilizes the finite volume method [83], and the computational domain includes both ferromagnetic and non-magnetic regions, as shown in Fig. 10(a) and (b). This method is compatible with unstructured meshes, and we therefore employed uniform unstructured meshes in the azimuthal direction to minimize any extraneous effects arising from the mesh geometry. The height of the nanotube in our µM simulation model is $700\,{\rm nm}$ (within the range of $z=-350$ to $+350\,{\rm nm}$). To minimize the impact of end effects on the DW, we conducted shift simulations with pulse width conditions designed to keep the DW within the range of $z=-200$ to $+200\,{\rm nm}$.
Figure 10: Top (a) and cross-sectional (b) views of the mesh structure in µM simulations. Green and pink represent the non-magnetic and magnetic regions, respectively.
Appendix B Derivation of magnetostatic interaction terms
In our PMA nanotube model, the surface and volume magnetic charge densities are given by
$\frac{\sigma_{{\rm m}}\left(\bm{x}\right)}{M_{\rm s}}$ =$=\begin{cases}-Q\tanh\frac{z-X}{\lambda}&{\rm at}\,r=R_{1}\\ +Q\tanh\frac{z-X}{\lambda}&{\rm at}\,r=R_{0}\\ \end{cases},$ (12)
$\frac{\rho_{\rm m}\left(\bm{x}\right)}{M_{\rm s}}$ $=\left(\frac{Q}{r}+\frac{\cos\phi}{\lambda}{\rm sech}{\frac{z-X}{\lambda}}\right){\tanh\frac{z-X}{\lambda}}.$ (13)
From Eqs. (2), (12), and (13), we obtain
$E_{{\rm dem,bb}}=$  $\,2\pi^{2}M_{{\rm s}}^{2}\lambda^{3}\int_{0}^{+\infty}{\rm d}\kappa\int_{\rho_{0}}^{\rho_{1}}{\rm d}\rho\,J_{0}\left(\rho\kappa\right)\rho\int_{\rho_{0}}^{\rho_{1}}{\rm d}\rho^{\prime}J_{0}\left(\rho^{\prime}\kappa\right)\rho^{\prime}\int_{-\infty}^{+\infty}{\rm d}x\int_{-\infty}^{+\infty}{\rm d}y$
×$\times\left(\frac{Q}{\rho}+\cos\phi\,{\rm sech}x\right)\left(\frac{Q}{\rho^{\prime}}+\cos\phi\,{\rm sech}y\right){\displaystyle\tanh x}\,{\displaystyle\tanh}y\,{\rm e}^{-\left|x-y\right|\kappa},$ (14)
$E_{{\rm dem,bs}}=$  $\,4\pi^{2}QM_{{\rm s}}^{2}\lambda^{3}\int_{0}^{+\infty}{\rm d}\kappa\int_{\rho_{0}}^{\rho_{1}}{\rm d}\rho\,J_{0}\left(\rho\kappa\right)\rho\left[J_{0}\left(\rho_{0}\kappa\right)\rho_{0}-J_{0}\left(\rho_{1}\kappa\right)\rho_{1}\right]\int_{-\infty}^{+\infty}{\rm d}x\int_{-\infty}^{+\infty}{\rm d}y$
×$\times\left(\frac{Q}{\rho}+\cos\phi\,{\rm sech}x\right){\displaystyle\tanh}x\,{\displaystyle\tanh}y\,{\rm e}^{-\left|x-y\right|\kappa},$ (15)
where $J_{n}\left(x\right)$ is the $n$-th order Bessel function of the first kind. The term $E_{{\rm dem,ss}}$, which is independent of $\phi$, is not included in the calculation as it does not affect the DW dynamics. Equations (14) and (15) include the terms proportional to $\cos\phi$, and Eq. (14) has the term proportional to $\cos\phi^{2}$. By dividing these by the tube cross-sectional area, the linear and quadratic terms of $\cos\phi$ can be expressed as
$w_{1}=$ $\,Q\frac{4\pi M_{{\rm s}}^{2}\lambda}{\rho_{1}^{2}-\rho_{0}^{2}}\cos\phi\int_{0}^{+\infty}{\rm d}\kappa\left[\rho_{1}J_{1}\left(\rho_{1}\kappa\right)-\rho_{0}J_{1}\left(\rho_{0}\kappa\right)\right]\int_{\rho_{0}}^{\rho_{1}}{\rm d}\rho\,J_{1}\left(\rho\kappa\right)\rho\int_{-\infty}^{+\infty}{\rm d}x\,{\displaystyle\tanh}x\int_{-\infty}^{+\infty}{\rm d}y\,\frac{{\tanh}y}{\cosh y}\,{\rm e}^{-\left|x-y\right|\kappa},$ (16)
$w_{2}=$ $\,\frac{2\pi M_{{\rm s}}^{2}\lambda}{\rho_{1}^{2}-\rho_{0}^{2}}\cos^{2}\phi\int_{0}^{+\infty}{\rm d}\kappa\left[\frac{\rho_{1}J_{1}\left(\rho_{1}\kappa\right)-\rho_{0}J_{1}\left(\rho_{0}\kappa\right)}{\kappa}\right]^{2}\int_{-\infty}^{+\infty}{\rm d}x\,\frac{{\tanh}x}{\cosh x}\int_{-\infty}^{+\infty}{\rm d}y\,\frac{{\tanh}y}{\cosh y}\,{\rm e}^{-\left|x-y\right|\kappa}.$ (17)
By comparing each result with Eqs. (3) and (4), we obtain Eqs. (6) and (7).